Contents

Thomas Achia: Generalized linear spatial modeling of HIV in Kenya 8 Nelson Afanador: Using R for Data Mining in Vaccine Manufacturing: Finding Needles in Biological

Haystacks 9 Ken Aho: Teaching statistics to biologists: the R-library asbio 10 Leslie AICHAOUI-DENEVE, Vincent FROMION, Matthieu JULES, Ludovic LE CHAT: Ba-

SyLiCA:A web interface for automatic process of Live Cell Array data using R 11 Andreas Alfons, Matthias Templ, Peter Filzmoser: The R package simFrame: An object-oriented ap-

proach towards simulation studies in statistics 12 Zack Almquist: US Census Spatial and Demographic Data in R: the UScensus2000-suite of packages 13 Lou Bajuk-Yorgan, Stephen Kaluzny: Making R accessible to Business Analysts with TIBCO Spotfire

14 Collin Bennett, Dave Locke, Robert Grossman, Steve Vejcik: Building Segmented Models Using R

and Hadoop 15 Kuntal Bhattacharyya, Pratim Datta, David Booth: To Do or Not To Do Business with a Country: A

Robust Classification Approach 16 Przemyslaw Biecek, Pawel Chudzian, Cezary Dendek, Justin Lindsey: Massively parallel analytics

for large datasets in R with nza package 17 Ralf Bierig, Jacek Gwizdka, Michael Cole, Nicholas Belkin: An Experiment Data Analysis Frame-

work: Evaluating Interactive Information Behaviour with R 18 Christopher Bilder, Boan Zhang, Frank Schaarschmidt, Joshua Tebbs: binGroup: A Package for

Group Testing 19 Richard Bilonick: Analysis of Data from Method Comparison Studies Using R, merror, and OpenMx 20 Carlin Brickner, Iordan Slavov, Rocco Napoli: Graphics Device Tabular Output 21 Christopher Brown: Venn : Powerful and High-Level and Set Manipulating ... a whole new way of

working with data 22 Giuseppe Bruno: Monte Carlo Simulation for Pricing European and American Basket option 23 Max Buot: Estimating a multivariate normal covariance matrix subject to a Loewner ordering 24 Gang Chen, Ziad Saad, Robert Cox: Statistical Analysis Programs in R for FMRI Data 25 Yauheniya Cherkas, Javier Cabrera, Birol Emir, Ha Nguyen: PfarMineR: An User-Friendly Expand-

able Front-End For Biopharmaceutical Applications with R 26 Karim Chine: User friendly distributed computing with R 27 Karim Chine: Elastic-R, a Google docs-like portal for data analysis in the Cloud 28 Christopher Brown: Eat your hashes! Hash come to R 29 Ettore Colombo, Gloria Ronzoni, Matteo Fontana: R role in Business Intelligence Software Architec-

ture 30 Drew Conway: Real-time network analysis in R using Twitter 31 Giles Crane, Cynthia Collins, Karin Mille: Evaluating Grant Applications with Generalized Chain

Block Designs in R 32 Karl-Dieter Crisman: Sage and R: Using R via the Sage notebook 33 Jeff Cromwell: Model Maker 1: Using Model Based Ontologies for Agent Based Estimation and Learning

of R Packages 34 Brian A. Danielak, Andrew Elby, Eric Kuo, Michael M. Hull: Using R to Assess Mathematical Sense-

Making in Introductory Physics Courses 35 Abhijit Dasgupta: Flexible report generation and literate programming using R and Python’s docutils

module 36 Neil Diamond: Using R for the Visualisation of Computer Experiments 37 Zubin Dowlaty, Deepak Bysani: Using R in an Event Driven Service Architecture 38 Adrian Dragulescu: Read, write, format Excel 2007 (xlsx) files 39 Dirk Eddelbuettel, Romain Francois: Rcpp: Seamless R and C++ integration 40 Dirk Eddelbuettel, Khanh Nguyen: RQuantLib: Bridging QuantLib and R 41 Steven Ellis: An Algorithm for Unconstrained Quadratically Penalized Convex Optimization 42 John Emerson, Michael Kane, Bryan Lewis: Real-time processing and analysis of data streams 43 Nadeem Faiz: A Simple Visualization of S and P 500 Index Performance 44 Michael Fay: Two-sided Exact Tests and Matching Confidence Intervals for Discrete Data 45 Ian Fellows: Deducer: A Graphical useR interface for everyone 46 Dai Feng, Luke Tierney: mritc - A package for MRI tissue classification 47 Michal Figurski, Leslie Shaw: Estimation of the Area-Under-the-Curve of Mycophenolic Acid using

population pharmacokinetic and multi-linear regression models simultaneously 48 Tom Filloon, Dave Dunlop: How to Effectively Visualize and Quantify a 3-D Image? 49 John Fox: The RcmdrPlugin.survival Package: A Survival-Analysis GUI for R 50 Romain Francois, Dirk Eddelbuettel: RProtoBuf: Protocol Buffers for R 51 Tal Galili: Blogging about R 52 Tal Galili, Shaul Shaul, Yoav Benjamini: Analyzing the Operational RNA Code for Amino Acids -

Using R 53 Christophe Genolini, Bernard Desgraupes, Lionel Riou Franca: R to LaTeX / HTML 54 Panayotis Giannakouros, Lihua Chen: Introducing computational thinking with free software in a math

for liberal arts course 55 Sheri Gilley: Designing a Flexible GUI for R 56 M. Helena Goncalves, M. Salome Cabral, Adelchi Azzalini: bild: a package for BInary Longitudinal

Data 57 Alex Guazzelli, Konstantinos Stathatos, Michael Zeller: PMML Execution of R Built Predictive So-

lutions 58 Ron Guida: Parallelizing a Computationally Intensive Financial R Application with Zircon Technology Christian Gunning: Using Rwave to detect synchrony of influenza between U.S. states 60 W. F. Guthrie, H. Liu: An Excel Interface for Functions in the meRology Package 61 Keith Halbert, Richard Heiberger, Erich Neuwirth: Export pivot table to R using RExcel 62 Michael Halter, Daniel R. Sisan, K.M. Mullen, Z.Q. John Lu: Statistical Analysis of Cell Population

Data 63 Bryan Hanson: ChemoSpec: An R Package for the Chemometric Analysis of Spectroscopic Data 64 Richard M. Heiberger, G. Jay Kerns: A Plot Method for "htest" Objects 65 Richard M. Heiberger: An Intermediate Course in Statistical Computing 66 Brian Hess, Michele Chambers: In-database analytics with R 67 Brian Hess, Michele Chambers: Analytics at Scale with R 68 Hsin-Ying Hsieh, Kun-Hsien Lin, Sun-Chong Wang: Cloud-R: toward a community-backed R in the

cloud 69 Mario Inchiosa, Cezary Dendek, Przemysaw Biecek: Scalable linear algebra with the nzmatrix package

70 John James: Use of and Using R as on Object Oriented Language 71 Landon Jensen: steReoscopy 72 Porzak Jim, Conway Drew, Pafka Szilard, Nash John: Panel on Starting and Building a Local R User

Group 73 Joseph Kahn, Peter Danenberg, Sourav Das, Derek Ayers: Model simulation and decision analysis

with the SimR package in R 74 Daniel Kaplan: Computers and the Teaching of Statistics 75 G. Jay Kerns: The IPSUR package: an Introduction to Probability and Statistics Using R 76 Rudiger Kessel: Automatic R-script generation for Monte Carlo Simulations 77 John Kloke, Joseph McKean, Patrick Kimes, Hilary Parker: Adaptive Nonparametric Statistics with

Applications to Gene Expression Data 78 Guido Knapp, Bimal Sinha, Dihua Xu: Extracting within-experiment precision of horticultural experi-

ments useful for meta-analysis 79 Ivan Kojadinovic, Jun Yan: Tests in Modeling Continuous Multivariate Distributions Using Copulas80 Fayaz Kondagula, Karl Molt: Infrared Spectrometric Purity Control of Chemical Substances using R81 Mieczyslaw Klopotek, Przemyslaw Biecek, Justin Lindsey: Simple Bayesian Networks on Netezza

Box 82 Max Kuhn: The caret Package: A Unified Interface for Predictive Models 83 Bill Ladd: Structured Text Access and Analysis 84 Andrew Lampitt: Placing the Power of R into Your Hands 85 Michael Larsen: Teaching Survey Sampling Theory using R 86 Autumn Laughbaum: Animated Graphics using R 87 Jonathan Lees: The Haiti Earthquake: Seismological Analysis Using R 88 Shengqiao Li, Donald Adjeroh, E. James Harner: Random KNN Classification and Regression 89 Hung-kung Liu, Steve Ellison, William Guthrie, Antonio Possolo: metRology - a new R package for

statistical metrology 90 Yingyun Liu: Using R for Data Simulation and Regression of Isothermal Titration Calorimetry of Proteins

with Alternative Conformations 91 Z.Q. John Lu, F. Potra, A. J. Wang: Fitting Multiphase Regression Models in R with Applications in

Microarray Experiments 92 Martin Maechler, Douglas Bates: Sparse Model Matrices for (Generalized) Linear Models 93 Ana Carolina CN Mafra, Ricardo Cordeiro, Luciana B Nucci, Celso Stephan: Generalized additive

models to nomial responses using bivariate Kernel: an solution to spatial analysis.Generalized additive models to nomial responses using bivariate Kernel: an solution to spatial analysis. 94

Patrick Mair, Reinhold Hatzinger: New developments for extended Rasch modeling in R 95 Frank Mannino, Richard Heiberger, Valerii Fedorov: Stochastic modeling and simulation in the design

of multicenter clinical trials 96 Marlene Marchena: The bullwhip effect under a generalized demand process: an R implementation 97 Inga Maslova: R package wfIMA: Wavelet-Functional Indexes of Magnetic Activity 98 Norman Matloff: Rdsm: Distributed (Quasi-)Threads Programming in R 99 Benjamin Mazzotta: Trade Cartograms: a Graphical Method for Dyadic Datasets 100 Muriel Mewissen, Thorsten Forster, Terry Sloan, Savvas Petrou: SPRINT: a Simple Parallel INTerface

to High Performance Computing and a Parallel R Function Library. 101 Margaret Mio-Julia, Arnau Mir, Monica J. Ruiz-Miro: R-TREE: Implementation of Decision Trees

using R 102 Arnau Mir, Margaret Miro-Julia, Monica J. Ruiz-Miro: Using R for Active Learning and Self-

assesment in an e-Learning Environment 103 Xuefei Mi, H. Friedrich Utz, Albrecht E. Melchinger: Integrated Development of the Software with

Literate Programming: An Example in Plant Breeding 104 Fausto Molinari, Martina Salvadori: R-Adamant: Financial technical analysis made easy 105 Junji Nakano, Ei-ji Nakama: Parallel Computing with R using GridRPC 106 Erich Neuwirth, Julia Theresa Csar: Prototyping Preventive Maintenance Tools with R 107 Erich Neuwirth: Investigating ODEs with R and spreadsheets 108 Pascal Neveu, Juliette Fabre: Using ontologies for R functions management 109 Derek Norton: David v. Goliath: How to build an R presence in a corporate SAS environment 110 Vincent Negre, Caroline Domerg, Juliette Fabre, Anne Tireau: Using R for data management in

ecophysiology Information Systems 111 Michael OConnell, Subra Subramanian: The use of R, S+ and Spotfire in the Oracle Life Sciences Data

Hub Environment 112 Jeroen Ooms: Web development with R 113 Anup Parikh, Kyle Covington: Red-R: A visual programming and data analysis framework 114 Pete Philipson, Ruwanthi Kolamunnage-Dona, Ines Sousa: Software for the joint modelling of longi-

tudinal and survival data: the JoineR package 115 Bill Pikounis, John Oleynick: The cg package for comparison of groups 116 Ashoka Polpitiya, Navdeep Jaitly, Konstantinos Petritis: A Tool for Quantitative Analysis of Pro-

teomics Data 117 Joerg Polzehl, Karsten Tabelow: Statistical Issues in Accessing Brain Functionality and Anatomy 118 Jim Porzak: Marketing Analytics in R 119 Setia Pramana, Dan Lin, Philippe Haldermans, Ziv Shkedy: IsoGeneGUI: a graphical user interface

for analyzing dose-response studies in microarray experiments 120 Richard Pugh, John James: Navigator: creating and managing complex graphics in a production envi-

ronment 121 Jeremy Raw: TravelR: Travel Demand Modeling in R 122 Andrew Redd: NppToR: R Interaction for Notepad++ 123 David Reinke, Steve Miller: R and BI – Integrating R with Open Source Business Intelligence Platforms

Pentaho and Jaspersoft 124 Jo Reynaerts, Ravi Varadhan, John C. Nash: The Convergence Properties of the BLP (1995) Contraction

Mapping and Alternative Algorithms in R 125 Yann Richet, David Ginsbourger, Olivier Roustant, Yves Deville: A Grid Computing Environment

for Design and Analysis of Computer Experiments 126 Bjorn Roelstraete, Yves Rosseel: spmR: An R package For fMRI Data Analysis Based On The SPM

Algorithms. 127 Gloria Ronzoni, Ettore Colombo, Matteo Fontana: R for Labour Market Policies 128 Dedi Rosadi: Teaching Time Series analysis course using RcmdrPlugin.Econometrics 129 Irina Roslyakova: Modified segmentation methods of quasi-stationary time series 130 Yves Rosseel: lavaan: an R package for structural equation modeling and more 131 Thomas Roth, Joachim Herrmann: Teaching Statistics in Quality Science using the R-Package quality-

Tools 132 Andrew Runnalls: CXXR and Add-on Packages 133 Paul Rustomji, Brent Henderson, Katie Mills, Bai Qifeng: R as a statistical engine for a water quality

trend analysis web -service 134 Jeffrey Ryan: Column Databases Made Easy with R 135 Jeffrey Ryan: Trading in Real Time with R and IBrokers 136 Sourish Saha, Vladimir Anisimov, Valerii Fedorov, Richard Heiberger: Drug Supply Modeling Soft-

ware 137 Daniel V. Samarov: Generalized Significance in Scale Space: The GS3 Package 138 Robert Samohyl, Elisa Henning: Forecast Monitoring via Multivariate Statistical Process Control with

R 139 Hector Sanz, Isaac Subirana, Joan Vila: Bivariate Analyses 140 Doug Schmidt: Adaptive Middleware and High Performance Software For Multi-core Deployments

Across Cloud Configurations 141 Neil Shah, Guruprasad Kora, Paul Breimyer, Yekaterina Shpanskaya: pR: Enabling Automatic Par-

allelization of Data-Parallel Tasks and Interfacing Parallel Computing Libraries in R with Application to Fusion Reaction Simulations 142

Yue Shentu: visualization of titrated dose and recurring events using R/ggplot2 143 Guria Sibnarayan: Diagnostics in Count Data Model 144 Jeff Skinner, VIvek Gopalan, Jason Barnett, Yentram Huyen: Automating biostatistics workflows for

bench scientists using R-based web-tools 145 David Smith: Evolving R for Use in Commercial Environments 146 Christopher Snyder, Keith Halbert: Plotting Advanced Mathematical Functions in Excel Using RExcel

147 Soeren Sonnenburg, Gunnar Raetsch, Sebastian Henschel, Christian Widmer: SHOGUN - A Large

Scale Machine Learning Toolbox 148 Ansgar Steland: A high-performance compiler for a subset of R 149 Ganesh Subramaniam, Ravi Varadhan, Simon Urbanek, Sam Epstein: tsX: An R package for the

exploratory analysis of a large collection of time-series 150 Eric Sun: Criss-Crossing the Org Chart: Predicting Colleague Interactions with R 151 Qiushan Tao: ChinaMap: Maps of China for analysing spatial data 152 Tom Taverner, Ashoka Polpitiya, Gordon A Anderson, Richard D Smith: RGtk2Extras and DanteR:

rapid GUI development for an "omics" R package 153 Paul Teetor: Analytics for Trading Financial Spreads 154 Matthias Templ, Karel Hron, Peter Filzmoser: robCompositions: An R-package for robust statistical

analysis of compositional data 155 Stefan Theussl, Kurt Hornik, David Meyer: Many Solvers, One Interface - ROI, the R Optimization

Infrastructure Package 156 Wesley Turner, Jeff Baumes, Phillipe Pebay, Thomas Otahal: Integration of R to VTK, Adding Statis-

tical Computation to a Visualization Toolkit 157 Heather Turner, David Firth: BradleyTerry2: Flexible Models for Paired Comparisons 158 Daniela Ushizima: Front propagation using fast marching in R 159 Mark van der Loo: Distribution based outlier detection with the extremevalues package 160 Ravi Varadhan: SQUAREM: An R package for Accelerating Slowly Convergent Fixed-Point Iterations

Including the EM and MM algorithms 161 Marie Vendettuoli, Dianne Cook, Heike Hofmann: Points, Curves and Haystacks: Datavis and

Metabolomics 162 John Verzani: The traitr package 163 Joan Vila, Montse Rue, Nuria Codern, Albert Sorribas: Teaching Statistics: An example of "How to”

improve the students’ statistical skills using individualized assignments 164 Hrishikesh Vinod: Time Series Inference Applications of R in Finance and Econometrics 165 Yue Wang, Narinder Nangia: Bayesian Monitoring of A Longitudinal Clinical Trial Using R2WinBUGS

166 Liang Wei, Brendan Kitts: Analyzing Direct Marketing Data with R 167 David Wheeler, Kai Yu, Melissa Friesen: Are there latent decision rules in expert occupational exposure

assessments? 168 Eric Wu, Patrick Mair, Peter Bentler: The REQS package for linking the SEM software EQS to R 169 Yihui Xie: Creating Animations with R 170 James Yen, Stephen Ellison: Analysis of interlaboratory studies using R and the metRology package 171 George Zhang: Social network analysis with R sna package 172 Jing Hua Zhao: Use of R in Genetic Epidemiology Designs 173 Alex Zolot: Generalized Linear Mixed Model with Spatial Covariates 174